OpenStudy (anonymous):

A wire of length 1.20m is fixed at one end and stretched. A transverse wave moves with a speed of 300 ms along the wire and is reflected at the fixed end. In the stationary wave set up, two successive nodes are seperated by a distance of 0.40m. What is the frequency for this mode of vibration?

OpenStudy (anonymous):

@ParthKohli @geerky42 @Vincent-Lyon.Fr

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

\[c=f lambda\] and they tell you that the nodes are a distance of the nodes are 0.40m apart and the distance between nodes is half a wavelength so lambda = 0.80m so you get \[c/0.80=f\]

OpenStudy (anonymous):

Thanks :) What are the lower frequencies obtainable?

OpenStudy (anonymous):

is it 125 Hz and 250Hz

OpenStudy (anonymous):

You should only get one frequency

OpenStudy (anonymous):

\[\frac{\lambda}{2} \times 1 = 1.20\] \[v = f_{0} \times \lambda\]

OpenStudy (anonymous):

this is the fundamental frequency right

OpenStudy (anonymous):

In this case it is not c but it is v. I used the formula for light in a vacuum

OpenStudy (anonymous):

it is V i know:0 so am i right?

OpenStudy (anonymous):

yes sorry for the confusion